AB INITIO AND DFT CALCULATIONS FOR THE VIBRATIONAL FREQUENCIES AND BARRIER TO PLANARITY OF CYCLOPENTENE AND ITS DEUTERATED ISOTOPOMERS

\underline{ABDULAZIZ AL-SAADI} and JAAN LAANE, Department of Chemistry, Texas A&M University, College Station, TX 77843-3255.

Ab initio and DFT calculations have been carried out for the cyclopentene molecule in order to analyze its structure and vibrational frequencies. The structure was calculated with MP2/6-311++G** and MP2/cc-pVTZ basis sets and these predicted puckering angles of 27.1^{∘} and 26.1^{∘}, respectively, as compared to the experimental far-infrared value of 26^{∘}. The barrier to planarity was calculated to be 247 cm^{−1}, slightly higher than the 233 cm^{−1} far-infrared value. The calculated vibrational frequencies from DFT-B3LYP/cc-pVTZ were compared to the experimental values for the d_{0}, d_{1}, d_{4}, and d_{8} isotopomers and several vibrational reassignments were made.